In the presence of excess OH-, the Al3+(aq) ion forms a hydroxide complex ion, Al(OH)4-. Calculate the concentration of free Al3+ ion when 1.32×10-2 mol Al(NO3)3(s) is added to 1.00 L of solution in which [OH- ] is held constant (buffered at pH 13.00). For Al(OH)4-, Kf = 1.1×1033.

[Al3+] = ?M
Answer #1

Al(NO₃)₂ + 40H - ALLOH) - at t=0 1.32x102 10t at t= leq (1.32x10 - ) (10-1.32x1024) & mole of Alat requires a moles of on- Cdel x 1033 = 1032x102 y x 4.96X106. 1032x102 4.96x100x1.1x1033 ly= 2.42 x 1025 m. [A1+3] = 2042 x 1029 m. so